OpenStudy (anonymous):

Solve A = 5(b + 2c) for c in terms of A and b. You must show all work.

OpenStudy (anonymous):

A = 5b + 10c => 10c = A-5b .. Now, solve for c..

OpenStudy (anonymous):

Ok so wold it be 10c - A/5=c

OpenStudy (anonymous):

Wait I mean c=A-5b/10